OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x^2 - 12 = -4

OpenStudy (anonymous):

is this really \[x^2-12=4\] or could it be \[x^2-12x=4\]?

OpenStudy (anonymous):

Oh my bad, the actual thing is x^2 - 12 =-4x